# The Common Data Format (CDF): A Standardized Format for Match-Data in Football (Soccer)

### v3

@@ 1 Introduction
Over the past two decades, professional sports organisations have increasingly relied on data-driven analysis to inform roster construction, tactical analysis, and marketing [Link2018, Lolli_2024]. This change has arisen for several reasons. First, there have been notable public examples of teams successfully employing data, which is most famously exemplified by baseball [Lewis2003]. Second, there has been a huge increase in the quantity and quality of data that can be collected, due in part to the growth of technology providers and associated financial incentives.
@@
record semantically meaningful “on-ball” events that occur during a match [Pappalardo2019a] such as passes, tackles and shots. Each event is annotated with information (e.g., location on the pitch, time-stamp, involved players). Events are typically collected manually by trained operators following a definition catalog. There are differences among vendors about the catalog of events and what information is recorded about each specific event.
@@
record the positions (i.e., x,y,z-coordinates) of all players, the ball, and the referees multiple times per second [Manafifard2017a]. This data is typically collected using an in-venue camera system or extracted from broadcast video footage with the help of bespoke image recognition software.
@@
The most powerful and meaningful analyses often require combining the above types of data [Rein2016]. Specifically, this is important in areas such as monitoring workload, performing tactical analysis, and scouting players.
Unfortunately, in practice, the way the data are collected and delivered pose a number of significant barriers that can prevent or limit performing such analyses. First, the definitions of key events are often ambiguous and vary across vendors, which is compounded by the fact that such events are often labeled by human annotators who make mistakes. Even in seemingly objective events such as shots, there are disagreements about how many of them occurred in a match [bialik:shot]. Second, the way data is delivered poses substantial challenges from an analysis perspective. Each source can use different file formats and structures. Moreover, they use different conventions about units of measure (e.g., local time vs. UTC, meters vs. yards), how the pitch is represented (e.g., coordinate system, playing direction), and how they refer to entities such as players and stadiums. Finally, the data sources are often collected independently of one another and it is highly non-trivial to align the data. On the one hand, it is necessary to cope with the same player being referred to using multiple naming conventions. On the other hand, time and/or location synchronizing of event and tracking data requires accounting for issues such as different clocks, different start times, and different coordinate systems [AnzerxG2021, VanRoy2023:ETSY]. Alas, addressing all these challenges requires a substantial investment in terms of both time and money. Consequently, data is often not exploited to the fullest extent possible.
@@ 2.3 Event Data
Event data or play-by-play data is a common type of data that breaks the game down into individual parts and is commonly collected in team sports such as football, ice hockey, rugby or basketball [Vracar2016]. This type of data records specific semantically meaningful events that occur during a match [Pappalardo2019a] and allows users including media outlets to query particular moments of the match. For example, relevant events in football include passes, tackles, and shots. Each event is annotated with information such as the players involved, the location on the pitch where the event occurred, the time of the event in the match, and the outcome of the event. Event data has traditionally been collected by companies such as Stats Perform [^1] Statsperform LLC, Chicago, https://www.statsperform.com/.; former AMISCO, Prozone and Opta., Sportec Solutions [^2] Sportec Solutions AG, Munich, Germany https://www.sportec-solutions.de/index.html., and Statsbomb [^3] Statsbomb Services Limited https://statsbomb.com/., who typically use human annotators to collect this data. Each company has their own catalog of events that they collect thus meaning datasets from different vendors do not always correspond. This is particularly true for events like tackles, interceptions or clearances where observational collection methods will lead to differences. With the launch of the FIFA Football Language [^4] FIFA, Zürich, https://www.fifatrainingcentre.com/en/resources-tools/football-language/ in 2022, a first framework has been offered to harmonise the terms and definitions used in football but more work is needed to ensure events are unambiguous. The development of automated event data detection from tracking data or video footage is another ongoing effort to ensure more uniform collection [FIFA-Event-Detection-2022].
@@ 2.4 Tracking Data
Tracking data record the positions of the players and the ball multiple times per second [Taberner2020]. The number of recordings per second is often referred to as the frame rate. Initial tracking systems recorded the location of a player (or referee) in 2D (i.e., x/y plane) based on the player’s center of mass, or center of trunk (in the case of upper mounted wearables). Henceforth this is referred to as center of body, as the player’s mass is rarely if ever calculated and thus is biomechanically inappropriate to use in this context. The position of the ball is often recorded with an additional z-coordinate that provides its height relative to the pitch surface. Newer approaches also record 3D positions for players, often augmented to include the location of various limbs (further referred to as landmarks) such as a player’s head, hands, feet, hips etc. This data is sometimes referred to as skeletal tracking data, or limb tracking data. Tracking data can be acquired in several different ways:
@@
Multiple studies have investigated the validity of specific tracking data systems (e.g., [Linke2020]). However, these tests cannot be conducted during official matches. Hence, it is difficult to fully quantify their accuracy.
@@ 2.6 Player Physical Data
Players also wear a number sensors during training sessions and matches to track their physical status. Most notably, the often wear a GNSS device that also contains an inertial measurement unit (e.g., accelerometer and gyroscope). Furthermore, it is possible that athletes wear other devices such as a heart rate monitor. These systems are used to derive a number of parameters such as the total distance covered and the number of accelerations. These are also typically subdivided by defining different intensity zones. A number of companies make these devices including Catapult Sports [^5] Catapult Sports, Melbourne, https://www.catapult.com/ and STATSports [^6] STATSports Group Limited, Ireland, https://pro.statsports.com/. Moreover, athletes are often asked to fill in questionnaires [Buchheit:2016] that might include information about their sleep or the rating of perceived exertion (RPE) [Borg1982] for a training session.
@@ 3.1 Data Specification
The set of events collected can greatly differ from one data provider to another. These differences are primarily due to three factors. The first factor is the tactical trends in the region where the data provider operates. For example, a data provider based in Europe might focus more on collecting events related to players pressuring the ball carrier, as pressing-based playing styles have gained traction in Europe. The second factor is the intended audience for the data feed. Data providers who cater to media companies and betting agencies have different priorities than those serving clubs and associations. The third factor is the complexity and the necessary time commitment required to collect certain types of events [bialik:annotate]. Data providers gather varying sets of events, but most tend to collect the most basic event types like shots and passes, which constitute the majority of on-ball events. However, since events like passes are not specified by the Laws of the Game [^8] IFAB, Laws of the Game, https://digitalhub.fifa.com/m/1cf301829f1cf996/original/ifab-laws-of-the-game-2020-21.pdf, data providers resort to their own definitions. Consequently, specific actions on the pitch may be classified differently by different providers, leading to inconsistencies in the resulting statistics and metrics [bialik:shot].
Furthermore, the breadth of the data specification also varies among data providers. Some providers limit their collection to actions performed by players, while others extend their scope to include meta events, such as changes in temperature or pitch conditions. Additionally, certain data providers incorporate performance metrics into their data specification, such as expected-goals values for shots [AnzerxG2021], expected pass values for passes [Anzer2022_xPass] or metrics on pressure [Andrienko2017], while others choose not to.
@@ 3.2 Data Representation
  • adopt different pitch coordinates. Some providers use normalized pitch coordinates, while others use actual pitch coordinates. The normalization of pitch coordinates introduces further challenges [Brandes-2023-Pitch], as pitches can vary in size, while most pitch markers are of fixed dimensions. For example, the penalty area and the penalty spot’s distance from the goal remain constant, regardless of the pitch dimensions.
@@ 4 Related Work
Although football data currently lacks a consensus and standard, there have been several individual attempts to simplify analysis by proposing unifying data representations. Interestingly, these are often backed by open-source software packages with permissive licenses (i.e., commercial use is allowed). The SPADL (Soccer Player Action Description Language) [Decroos2019] proposes a tabular representation for a subset of the key information recorded in event data. This puts the data in a natural format and facilitates computing advanced performance metrics from event data such as expected goals (xG) [robberechts2020:xg] and Expected Threat (xT) [rudd2011framework, Singh:xT]. The software package supports parsing data from several different data providers and this package is used in both research and industry.
For tracking data, FIFA and F.C. Barcelona jointly created global standards for Electronic Performance and Tracking Systems (EPTS) [fifa:epts]. More recently, there are open-source packages such as kloppy [kloppy] and floodlight [raabe2022floodlight] that support parsing both tracking and event data. Moreover, there are several open-source algorithms for synchronizing tracking and event data [oonk:databallpy, VanRoy2023:ETSY].
While these initiatives are acknowledged by data providers and widely used by researchers and practitioners, an effort to harmonize all match data, supported by a large group of stakeholders has not yet culminated. Similar to the work of Robertson et al. [robertson2023development], the CDF aims to provide a framework for handling football data that offers possible solutions to common challenges as outlined in the previous section.
@@ 5.1 Specification of the CDF Schema
The CDF specifies a minimal schema for five types of match data: match sheet data, video footage, event data, tracking data, and match meta data. However, it is not mandatory that all five types of data are available. The only two requirements on availability are that (1) match sheet data must be included, and (2) if either event or tracking data is provided, then match meta data must also be included. Additionally, if both event and tracking data are provided, these data sources should be synchronized in time and location (see e.g., [oonk:databallpy, VanRoy2023:ETSY, AnzerxG2021, Anzer2022_xPass]).
@@
Table 3 provides the schema for event data. The CDF only includes events that are either objective or whereby clear consensus exists on their definition (irrespective of the collection method): Events that do not meet this threshold are excluded from the CDF. The events that meet this requirement are listed in event_(sub)type fields of the provided schema. [^11] There is ongoing work from a FIFA expert group that is developing a touched-based event model [FIFA_2023_Whitepaper].
@@ 5.2 Representational Conventions in the CDF
An advantage of a coordinate system with both positive and negative values is that it allows for custom normalization simply by multiplying various coordinates by $-1$.
@@ 5.3 Structure of Delivered Data
We have selected these formats because they confer three important benefits. First, they are relatively compact in terms of memory used per file (compared to other commonly used file formats like JSON or XML), and are also efficient in terms of maximum memory capacity used when loading or streaming into memory. Second, a well-structured (and versioned) JSON (Lines) file format will allow providers to enrich their data with more features as time progresses and new metrics become available without breaking any existing infrastructure. Third, all popular data processing tools ($pandas$, $PySpark$, $polars$ and $R$) support this file type, and therefore further post-processing can be left to each individual receiver.
@@ 5.3.1 Meta- and Match Sheet Data File Format
The match sheet and meta data should be delivered in a JSON file (.$json$). Figure 2 shows a simplified example of match sheet data where some mandatory CDF fields are omitted for readability. The optional fields are shown in Table 7. Figure 3 depicts a simplified example of meta data that does not shown all mandatory CDF fields for readability. The optional fields are shown in Table 11.
Finally, if provided, skeletal tracking data should be accompanied by a skeletal hierarchy in the meta data under ”meta/landmarks_nodes”. It should adhere to the format shown in Figure 6 which shows a simplified hierarchy of a standardized T-pose. The format is borrowed from the glTF2.0 file format  [gltf2_specification]. The ”name” field specifies the joint under consideration, it follows a snake_case formatting with the ”right”, ”left” indicators as a suffix. The ”rotation” field is a quaternion defining the rotation of the joint in a T-pose, this can generally be left as [0.0, 0.0, 0.0, 1.0], but should be adjusted for joints that do not make right angles with their parent node (e.g. ”thumb_right” or ”eye_left”). Translation should follow the ”Y-up” approach such that $x$, $y$ and $z$ components specify the orientation in relation to their parent node. For example ”hip” has a child-node ”spine” (as defined by the landmark node indices in ”children”) and there for [0, 1, 0] indicates the spine goes up (1) in the Y-axis and does not deviate in the $x$ or $z$ direction compared to the hip.
@@
Figure 3: Example of a non-nested JSON (.$json$) meta data file with stadium, player, and team information.
@@ 5.3.2 Event and Tracking Data File Format
Figure 4 illustrates the first two events of the first half. Figures 5 shows a simplified example for two frames of player center of body tracking data. Note that we are only showing the mandatory CDF fields. These can be augmented with the optional fields (see Tables 9 and 10). If skeletal (limb) tracking is provided, it should follow the same structure as the tracking data JSON Lines format. Instead of a players $x$ and $y$ coordinates, the provider would provide coordinates for each landmark (e.g., $left\_hip\_x$, $right\_hip\_y$, $right\_hip\_x$) as specified by Table 5. The coordinate system should be the same used in the center of body tracking data, e.g., $x$, $y$, $z$ and are relative to the pitch. The field $is\_visible$ indicates if the part was actually measured, or inferred via an algorithm due to being hidden. Providers have flexibility over which body parts they track. This should be communicated beforehand.
@@
Figure 4: Example of JSON Lines (.$jsonl$) that covers the mandatory CDF fields for Event Data. The file describes two events which are expanded over multiple lines for readability.
@@
Figure 5: Example of the JSON Lines (.$jsonl$) for the Center of Body Tracking Data representation. Two frames are shown and they are expanded over multiple lines for readability. We omit the positions of all players for readability.

### v4

@@ 1 Introduction
Over the past two decades, professional sports organisations have increasingly relied on data-driven analysis to inform roster construction, tactical analysis, and marketing [17, 19]. This change has arisen for several reasons. First, there have been notable public examples of teams successfully employing data, which is most famously exemplified by baseball [16]. Second, there has been a huge increase in the quantity and quality of data that can be collected, due in part to the growth of technology providers and associated financial incentives.
@@
record semantically meaningful “on-ball” events that occur during a match [22] such as passes, tackles and shots. Each event is annotated with information (e.g., location on the pitch, time-stamp, involved players). Events are typically collected manually by trained operators following a definition catalog. There are differences among vendors about the catalog of events and what information is recorded about each specific event.
@@
record the positions (i.e., x,y,z-coordinates) of all players, the ball, and the referees multiple times per second [20]. This data is typically collected using an in-venue camera system or extracted from broadcast video footage with the help of bespoke image recognition software.
@@
The most powerful and meaningful analyses often require combining the above types of data [24]. Specifically, this is important in areas such as monitoring workload, performing tactical analysis, and scouting players.
Unfortunately, in practice, the way the data are collected and delivered pose a number of significant barriers that can prevent or limit performing such analyses. First, the definitions of key events are often ambiguous and vary across vendors, which is compounded by the fact that such events are often labeled by human annotators who make mistakes. Even in seemingly objective events such as shots, there are disagreements about how many of them occurred in a match [5]. Second, the way data is delivered poses substantial challenges from an analysis perspective. Each source can use different file formats and structures. Moreover, they use different conventions about units of measure (e.g., local time vs. UTC, meters vs. yards), how the pitch is represented (e.g., coordinate system, playing direction), and how they refer to entities such as players and stadiums. Finally, the data sources are often collected independently of one another and it is highly non-trivial to align the data. On the one hand, it is necessary to cope with the same player being referred to using multiple naming conventions. On the other hand, time and/or location synchronizing of event and tracking data requires accounting for issues such as different clocks, different start times, and different coordinate systems [2, 30]. Alas, addressing all these challenges requires a substantial investment in terms of both time and money. Consequently, data is often not exploited to the fullest extent possible.
@@ 2.3 Event Data
Event data or play-by-play data is a common type of data that breaks the game down into individual parts and is commonly collected in team sports such as football, ice hockey, rugby or basketball [33]. This type of data records specific semantically meaningful events that occur during a match [22] and allows users including media outlets to query particular moments of the match. For example, relevant events in football include passes, tackles, and shots. Each event is annotated with information such as the players involved, the location on the pitch where the event occurred, the time of the event in the match, and the outcome of the event. Event data has traditionally been collected by companies such as Stats Perform [^1] Statsperform LLC, Chicago, https://www.statsperform.com/.; former AMISCO, Prozone and Opta., Sportec Solutions [^2] Sportec Solutions AG, Munich, Germany https://www.sportec-solutions.de/index.html., and Statsbomb [^3] Statsbomb Services Limited https://statsbomb.com/., who typically use human annotators to collect this data. Each company has their own catalog of events that they collect thus meaning datasets from different vendors do not always correspond. This is particularly true for events like tackles, interceptions or clearances where observational collection methods will lead to differences. With the launch of the FIFA Football Language [^4] FIFA, Zürich, https://www.fifatrainingcentre.com/en/resources-tools/football-language/ in 2022, a first framework has been offered to harmonise the terms and definitions used in football but more work is needed to ensure events are unambiguous. The development of automated event data detection from tracking data or video footage is another ongoing effort to ensure more uniform collection [31].
@@ 2.4 Tracking Data
Tracking data record the positions of the players and the ball multiple times per second [29]. The number of recordings per second is often referred to as the frame rate. Initial tracking systems recorded the location of a player (or referee) in 2D (i.e., x/y plane) based on the player’s center of mass, or center of trunk (in the case of upper mounted wearables). Henceforth this is referred to as center of body, as the player’s mass is rarely if ever calculated and thus is biomechanically inappropriate to use in this context. The position of the ball is often recorded with an additional z-coordinate that provides its height relative to the pitch surface. Newer approaches also record 3D positions for players, often augmented to include the location of various limbs (further referred to as landmarks) such as a player’s head, hands, feet, hips etc. This data is sometimes referred to as skeletal tracking data, or limb tracking data. Tracking data can be acquired in several different ways:
@@
Multiple studies have investigated the validity of specific tracking data systems (e.g., [18]). However, these tests cannot be conducted during official matches. Hence, it is difficult to fully quantify their accuracy.
@@ 2.6 Player Physical Data
Players also wear a number sensors during training sessions and matches to track their physical status. Most notably, the often wear a GNSS device that also contains an inertial measurement unit (e.g., accelerometer and gyroscope). Furthermore, it is possible that athletes wear other devices such as a heart rate monitor. These systems are used to derive a number of parameters such as the total distance covered and the number of accelerations. These are also typically subdivided by defining different intensity zones. A number of companies make these devices including Catapult Sports [^5] Catapult Sports, Melbourne, https://www.catapult.com/ and STATSports [^6] STATSports Group Limited, Ireland, https://pro.statsports.com/. Moreover, athletes are often asked to fill in questionnaires [11] that might include information about their sleep or the rating of perceived exertion (RPE) [9] for a training session.
@@ 3.1 Data Specification
The set of events collected can greatly differ from one data provider to another. These differences are primarily due to three factors. The first factor is the tactical trends in the region where the data provider operates. For example, a data provider based in Europe might focus more on collecting events related to players pressuring the ball carrier, as pressing-based playing styles have gained traction in Europe. The second factor is the intended audience for the data feed. Data providers who cater to media companies and betting agencies have different priorities than those serving clubs and associations. The third factor is the complexity and the necessary time commitment required to collect certain types of events [6]. Data providers gather varying sets of events, but most tend to collect the most basic event types like shots and passes, which constitute the majority of on-ball events. However, since events like passes are not specified by the Laws of the Game [^8] IFAB, Laws of the Game, https://digitalhub.fifa.com/m/1cf301829f1cf996/original/ifab-laws-of-the-game-2020-21.pdf, data providers resort to their own definitions. Consequently, specific actions on the pitch may be classified differently by different providers, leading to inconsistencies in the resulting statistics and metrics [5].
Furthermore, the breadth of the data specification also varies among data providers. Some providers limit their collection to actions performed by players, while others extend their scope to include meta events, such as changes in temperature or pitch conditions. Additionally, certain data providers incorporate performance metrics into their data specification, such as expected-goals values for shots [2], expected pass values for passes [3] or metrics on pressure [1], while others choose not to.
@@ 3.2 Data Representation
  • adopt different pitch coordinates. Some providers use normalized pitch coordinates, while others use actual pitch coordinates. The normalization of pitch coordinates introduces further challenges [10], as pitches can vary in size, while most pitch markers are of fixed dimensions. For example, the penalty area and the penalty spot’s distance from the goal remain constant, regardless of the pitch dimensions.
@@ 4 Related Work
Although football data currently lacks a consensus and standard, there have been several individual attempts to simplify analysis by proposing unifying data representations. Interestingly, these are often backed by open-source software packages with permissive licenses (i.e., commercial use is allowed). The SPADL (Soccer Player Action Description Language) [12] proposes a tabular representation for a subset of the key information recorded in event data. This puts the data in a natural format and facilitates computing advanced performance metrics from event data such as expected goals (xG) [25] and Expected Threat (xT) [27, 28]. The software package supports parsing data from several different data providers and this package is used in both research and industry.
For tracking data, FIFA and F.C. Barcelona jointly created global standards for Electronic Performance and Tracking Systems (EPTS) [13]. More recently, there are open-source packages such as kloppy [32] and floodlight [23] that support parsing both tracking and event data. Moreover, there are several open-source algorithms for synchronizing tracking and event data [21, 30].
While these initiatives are acknowledged by data providers and widely used by researchers and practitioners, an effort to harmonize all match data, supported by a large group of stakeholders has not yet culminated. Similar to the work of Robertson et al. [26], the CDF aims to provide a framework for handling football data that offers possible solutions to common challenges as outlined in the previous section.
@@ 5.1 Specification of the CDF Schema
The CDF specifies a minimal schema for five types of match data: match sheet data, video footage, event data, tracking data, and match meta data. However, it is not mandatory that all five types of data are available. The only two requirements on availability are that (1) match sheet data must be included, and (2) if either event or tracking data is provided, then match meta data must also be included. Additionally, if both event and tracking data are provided, these data sources should be synchronized in time and location (see e.g., [21, 30, 2, 3]).
@@
Table 3 provides the schema for event data. The CDF only includes events that are either objective or whereby clear consensus exists on their definition (irrespective of the collection method): Events that do not meet this threshold are excluded from the CDF. The events that meet this requirement are listed in event_(sub)type fields of the provided schema. [^11] There is ongoing work from a FIFA expert group that is developing a touched-based event model [14].
@@ 5.2 Representational Conventions in the CDF
An advantage of a coordinate system with both positive and negative values is that it allows for custom normalization simply by multiplying various coordinates by 1-1.
@@ 5.3 Structure of Delivered Data
We have selected these formats because they confer three important benefits. First, they are relatively compact in terms of memory used per file (compared to other commonly used file formats like JSON or XML), and are also efficient in terms of maximum memory capacity used when loading or streaming into memory. Second, a well-structured (and versioned) JSON (Lines) file format will allow providers to enrich their data with more features as time progresses and new metrics become available without breaking any existing infrastructure. Third, all popular data processing tools (pandaspandas, PySparkPySpark, polarspolars and $R$) support this file type, and therefore further post-processing can be left to each individual receiver.
@@ 5.3.1 Meta- and Match Sheet Data File Format
The match sheet and meta data should be delivered in a JSON file (.json.json). Figure 2 shows a simplified example of match sheet data where some mandatory CDF fields are omitted for readability. The optional fields are shown in Table 7. Figure 3 depicts a simplified example of meta data that does not shown all mandatory CDF fields for readability. The optional fields are shown in Table 11.
Finally, if provided, skeletal tracking data should be accompanied by a skeletal hierarchy in the meta data under ”meta/landmarks_nodes”. It should adhere to the format shown in Figure 6 which shows a simplified hierarchy of a standardized T-pose. The format is borrowed from the glTF2.0 file format  [15]. The ”name” field specifies the joint under consideration, it follows a snake_case formatting with the ”right”, ”left” indicators as a suffix. The ”rotation” field is a quaternion defining the rotation of the joint in a T-pose, this can generally be left as [0.0, 0.0, 0.0, 1.0], but should be adjusted for joints that do not make right angles with their parent node (e.g. ”thumb_right” or ”eye_left”). Translation should follow the ”Y-up” approach such that $x$, $y$ and $z$ components specify the orientation in relation to their parent node. For example ”hip” has a child-node ”spine” (as defined by the landmark node indices in ”children”) and there for [0, 1, 0] indicates the spine goes up (1) in the Y-axis and does not deviate in the $x$ or $z$ direction compared to the hip.
@@
Figure 3: Example of a non-nested JSON (.json.json) meta data file with stadium, player, and team information.
@@ 5.3.2 Event and Tracking Data File Format
Figure 4 illustrates the first two events of the first half. Figures 5 shows a simplified example for two frames of player center of body tracking data. Note that we are only showing the mandatory CDF fields. These can be augmented with the optional fields (see Tables 9 and 10). If skeletal (limb) tracking is provided, it should follow the same structure as the tracking data JSON Lines format. Instead of a players $x$ and $y$ coordinates, the provider would provide coordinates for each landmark (e.g., left_hip_xleft\_hip\_x, right_hip_yright\_hip\_y, right_hip_xright\_hip\_x) as specified by Table 5. The coordinate system should be the same used in the center of body tracking data, e.g., $x$, $y$, $z$ and are relative to the pitch. The field is_visibleis\_visible indicates if the part was actually measured, or inferred via an algorithm due to being hidden. Providers have flexibility over which body parts they track. This should be communicated beforehand.
@@
Figure 4: Example of JSON Lines (.jsonl.jsonl) that covers the mandatory CDF fields for Event Data. The file describes two events which are expanded over multiple lines for readability.
@@
Figure 5: Example of the JSON Lines (.jsonl.jsonl) for the Center of Body Tracking Data representation. Two frames are shown and they are expanded over multiple lines for readability. We omit the positions of all players for readability.